When PCl5 is heated it gasifies and dissociates into PCl3 and Cl2. Thedensity of the gas mixture at 200°C is 70.2. Find the degree of dissociation of PCl5 at 200°C.

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
initial vapour densuty/vapour density at equilibrium=mole at equilibrium/mole initialPCL5 -----PCL3 +CL2let the initial concentration be Cconcentration of PCL3=C-Cx,x is the degree of dissociationPCL3=CxCL2=Cxtotal mole at equilibrium=C +Cxinitial mole=CthereforeC +Cx/C =104.25/70.21+x=1.485x=.485  RegardsArun (askIITians forum expert)